# Regularizing Extrapolation in Causal Inference

### v1

@@ 3.1 Characterizing asymmetry-induced bias
Thus far we have presented a conservative nonparametric bound. We now provide a slightly refined characterization by noting that the extent of the bias induced by negative weights is driven by the asymmetry in $\mu$. We do so by considering the decomposition of the $\mu$ into its even and odd components, i.e., $\mu(x)=\mu_{e}(x)+\mu_{o}(x)$. By the definition of odd functions, we have $\mu_{o}(x)=\mu_{o}(-x)$, so we can bound the risk by bounding the worst-case risk of $\hat{\boldsymbol{w}}$ using the assumed Hölder constants $a$ and $\alpha$ and isolating the effect of the even component. The formal statement is given below in Proposition 3.1 the proof is given in Appendix B.
@@
Let $\hat{\mu}(x^{*})=\sum_{i=1}^{n}\hat{w}_{i}Y_{i}$ be the estimate of $\mu(x^{*})$ with weights estimated via Equation 5. Given $Y_{i}=\mu(X_{i})+\epsilon_{i}$ where $\epsilon_{i}$ are independent random variables with $\mathbb{E}[\epsilon_{i}]=0$ and finite second moment $\sigma^{2}=\mathbb{E}[\epsilon_{i}^{2}]$, and $\mu$ is Hölder continuous with constants $a$ and $\alpha$. If $\epsilon_{i}$ are sub-Gaussian [^1] We assume mean zero sub-Gaussian noise, analogous results can be obtained with this assumption replaced by bounded noise. with parameter $\sigma$, then with probability at least $1-\delta$,
@@
$B_{\text{even}}(x^{*})=\left|\sum_{i=1}^{n}\hat{w}_{i}[\mu_{e}(X_{i})-\mu_{e}(x^{*})]+2\sum_{i=1}^{n}|\hat{w}_{i}|\mathbf{1}(\hat{w}_{i}<0)a\|X_{i}-x^{*}\|^{\alpha}\right|$ (4)
@@
The worst-case bound provided earlier is recovered if $\mu$ is an even function, i.e., contains no odd component. One of the fundamental limitations of Proposition 3.1 is that it requires access to the separate even and odd functions constituting $\mu$ which are latent in practice. In our proposed estimator and empirical application we address this by preferring the conservative form which assumes the worst case form. For completeness we all provide the following proposition that provides an empirical analog of Proposition 3.1. Intuitively, when the observations of $X$ are symmetric, i.e., for every $X_{i}$ $X_{i}$ is also in the dataset, then we can recover both the even and odd functions. In practice, because this symmetry is unlikely to hold we approximate it with a one-nearest neighbor and incorporate the induced uncertainty into the bound using the Hölder constants. The formal statement is below, the proof is deferred to Appendix B.
@@
Let $\hat{\mu}(x^{*})=\sum_{i=1}^{n}\hat{w}_{i}Y_{i}$ be the estimate of $\mu(x^{*})$ with weights estimated via Equation (5). Given $Y_{i}=\mu(X_{i})+\epsilon_{i}$ where $\epsilon_{i}$ are independent sub-Gaussian random variables with parameter $\sigma$, and $\mu$ is Hölder continuous with constants $a$ and $\alpha$. Let $I_{\text{paired}}=\{i:-X_{i}\in\{X_{1},\ldots,X_{n}\}\}$, $I_{\text{nn}}=\{1,\ldots,n\}\setminus I_{\text{paired}}$, and for each $i\in I_{\text{nn}}$, define $j^{*}(i)=\arg\min_{j\neq i}\|X_{j}-(-X_{i})\|$. Then with probability at least $1-\delta$:
$|\mu(x^{*})-\hat{\mu}(x^{*})|\leq\left|\sum_{i\in I_{\text{paired}}}\hat{w}_{i}\frac{Y_{i}+Y_{-i}}{2}+\sum_{i\in I_{\text{nn}}}\hat{w}_{i}\frac{Y_{i}+Y_{j^{*}(i)}}{2}-\mu_{e}(x^{*})\sum_{i=1}^{n}\hat{w}_{i}\right|$
$\quad+\sum_{i\in I_{\text{nn}}}|\hat{w}_{i}|a\|X_{j^{*}(i)}-(-X_{i})\|^{\alpha}+2\sum_{i=1}^{n}|\hat{w}_{i}|\mathbf{1}(\hat{w}_{i}<0)a\|X_{i}-x^{*}\|^{\alpha}$
$\quad+\sigma\|\hat{w}\|_{2}\sqrt{2\log(6/\delta)}+\frac{\sigma}{\sqrt{2}}(\left\|\hat{w}_{I_{\text{paired}}}\right\|_{2}+\left\|\hat{w}_{I_{\text{nn}}}\right\|_{2})\sqrt{2\log(6/\delta)}$
where $\mu_{e}(x^{*})$ is bounded using the closest observation $j^{*}=\arg\min_{j=1,\ldots,n}\|X_{j}-x^{*}\|$,
$|\mu_{e}(x^{*})|\leq\left|\frac{Y_{j^{*}}+Y_{j^{*}(j^{*})}}{2}\right|+\sigma\sqrt{2\log(6/\delta)}+a\|X_{j^{*}(j^{*})}-(-X_{j^{*}})\|^{\alpha}+a\|X_{j^{*}}-x^{*}\|^{\alpha}$
Proposition 3.2 can be used as a companion to Chattopadhyay and Zubizarreta (2023), who propose computing the effective sample size of units with negative weight, $\frac{\sum_{i}\mathbf{1}[w_{i}<0]|w_{i}|}{\sum_{i}|w_{i}|}$, to indicate the extent to which the estimate relies on extrapolation and parametric assumptions.
@@ 3.2 Learning Estimator
We now propose a estimator to learn weights $\boldsymbol{w}$} that directly control the error bound in Equation (2). To do so, we modify the standard balancing weights optimization problem in Equation (1) by using the Lagrangian form of the non-negative restriction, rather than the hard constraint. Thus, the combined estimator minimizes the error bound by controlling three terms: covariate imbalance, dispersion of the weights, and level of extrapolation:
$\hat{\mathbf{w}}\in\arg\min_{\mathbf{w}}\underbrace{\|\sum_{i}^{n}w_{i}{\boldsymbol{X}}_{i}-{\boldsymbol{x}}^{\star}\|^{2}_{p}}_{(a)}+\lambda\underbrace{\|\mathbf{w}\|^{2}_{2}}_{(b)}+\gamma\underbrace{\left\|\mathbbm{1}(w_{i}<0)|w_{i}|\left(\|{\boldsymbol{X}}_{i}\|^{\alpha}_{2}\right)\right\|_{p}}_{(c)}.$ (5)
Where:
@@
The standard balancing weights problem in Equation (1) only focuses on a single bias-variance tradeoff: trading off covariate imbalance in the first term — which directly introduces bias — and the norm of the weights in the second term — which directly controls the estimator variance. By contrast, the new optimization problem (5) has a more elaborate bias-bias-variance trade-off. Allowing for negative weights introduces an additional trade-off between the first two terms and term (c): when the target unit lies outside the convex hull of the training points, controlling imbalance often requires some $w_{i}$ values to be negative, which also increases the norm $\|{\boldsymbol{w}}\|_{2}$. Allowing negative weights also necessitates reliance on parametric assumptions for extrapolation.
@@
Term (c) mitigates the risk of biased estimation by regulating the contribution of negative weights. For $\gamma=0$, Equation (5) recovers a standard, unconstrained balancing weights problem as in Equation (1). At the other extreme $\gamma\to\infty$ is equivalent to a hard non-negativity constraint. Increasing $\gamma$ constrains extrapolation, reducing bias due to possible violations of parametric assumptions and limiting $\|{\boldsymbol{w}}\|_{2}$, but worsening bias due to insufficient balance in term (a).
4 Synthetic Data Study
We evaluate our approach using synthetic data with both (i) linear and (ii) nonlinear data-generating processes (DGPs) where the target point lies outside the convex hull of training points, creating a challenging extrapolation scenario with limited sample size (10 training units) (see Figure 6). We provide a brief summary in the main text and defer the full description and analysis to Appendix C.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 1: Results on synthetic data generated using linear DGP. (a) Estimation error measured as mean squared error, (b) balance between the weighted source and target populations, (c) extent of extrapolation measured as negative influence – contribution on units with negative weights, (d) L2 norm $\mathbf{w}$ capturing asymptotic variance.
For the linear setting where parametric assumptions hold, Figure 7(a) shows that estimation error increases as we regularize extrapolation (as expected). This is in congruence lack of balance shown in Figure 7(b). As the parametric assumption holds, relying on it for extrapolation yields optimal estimates. However, for the nonlinear DGP where parametric assumptions are violated through quadratic and interaction terms, this is not true. Figure 8 illustrates our theoretical bias-bias-variance tradeoff. Small amounts of extrapolation remain beneficial due to the linear component, but excessive extrapolation leads to high error rates due to assumption violations. Regularization reduces extrapolation bias from parametric misspecification but increases distributional imbalance bias.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 2: Results on synthetic data generated using non-linear DGP. (a) Estimation error measrued as mean squared error, (b) balance between the weighted source and target populations, (c) extent of extrapolation measured as negative influence, (d) L2 norm $\mathbf{w}$ capturing asymptotic variance.
5 Application: Generalizing Opioid Use Disorder Trial Evidence
We demonstrate our framework using data from the START trial (Saxon et al., 2013), which compared buprenorphine versus methadone for treating opioid use disorder.
Data Description.
@@ 5 Application: Generalizing Opioid Use Disorder Trial Evidence
The Starting Treatment With Agonist Replacement Therapies (START) trial, initiated in 2006, was a multi-center study comparing buprenorphine versus methadone in treating opioid use disorder (Saxon et al., 2013; Hser et al., 2014). The trial enrolled 1,271 participants, who were randomized in a 2:1 ratio to receive either buprenorphine or methadone. Methadone was found to have higher rates of patient retention in treatment compared to buprenorphine (though buprenorphine in this trial was given in an unusual way to mimic methadone medication administration—requiring near daily clinic visits of participants) (Hser et al., 2014). Our analysis focuses on the outcome of relapse to regular opioid use within 24 weeks of medication assignment, defined as non-study opioid use for four consecutive weeks or daily use for seven consecutive days. Data on opioid use were collected through urine drug screens and self-reports, with relapse assessment beginning 20 days post-randomization to account for residual drug presence during stabilization.
@@
Parikh et al. (2025) identified that Latina women with a pre-treatment history of amphetamine and benzodiazepine use were underrepresented in the START trial relative to the target population, highlighting a practical violation of the positivity assumption. In this study, we estimate the average treatment effects (ATE) for this underrepresented subgroup using our proposed framework alongside standard linear regression and inverse probability weighting (IPW) estimators. We also evaluate the sensitivity of these estimates to parametric assumptions.
@@
The target sample is drawn from the 2015–2017 Treatment Episode Dataset - Admissions (TEDS-A), which includes data on individuals entering publicly funded substance use treatment programs across 48 states (excluding Oregon and Georgia) and the District of Columbia. Our analysis focuses on Latina women with a pre-treatment history of amphetamine and benzodiazepine use.
Analysis.
@@
Using linear regression, the estimated treatment effect for our subgroup of interest is 0.278-0.278, indicating that relapse rates are approximately 28 percentage points lower under methadone compared to buprenorphine. In contrast, the IPW estimate is 0.014-0.014, suggesting that both treatments are similarly effective. However, as shown in Figure 11, the negative influence of the linear regression-based estimate is 35% and 40% for $T=0$ and $T=1$, respectively, compared to 0% for IPW. Despite this, Figure 10 demonstrates that the IPW estimator achieves worse covariate balance than linear regression. These findings reveal that while linear regression relies heavily on parametric assumptions and negative weights, it may be biased if these assumptions are violated. Conversely, the IPW estimator avoids additional parametric assumptions but introduces bias due to poor covariate balance and violations of the positivity assumption.
Figure 3: Average Treatment Effects for the Target Sample from TEDS-A of Hispanic Females who have a history of Amphetamize and Benzodiazepine use. Each hue corresponds to a value of $\lambda$ and the x-axis corresponds to different values of $\gamma$ (on log scale). The dashed line represents the estimate using linear regression and the dotted represents the estimate using inverse probability weighting (IPW).
We apply our proposed framework to address these issues, which regularizes extrapolation to mitigate reliance on extreme weights. By varying $\gamma$ from 0.01 to 10, we examine how treatment effect estimates shift with increasing regularization of negative weights. Without regularization, the estimates converge with those from linear regression. However, as regularization intensifies, the estimates smoothly shift towards zero and occasionally change the sign from negative to positive for smaller values of $\lambda$. This sensitivity underscores the influence of assumptions on the point estimates. While increasing $\gamma$ reduces negative influence (Figure 11), it worsens covariate balance, as reflected in higher RMSE values (Figure 10). Thus, our framework highlights a trade-off between minimizing reliance on parametric assumptions and achieving optimal covariate balance.
Figure 4: Balance between the trial and the target samples measured as the root mean squared error (RMSE) for different values of $\gamma$ and $\lambda$ along with implied linear regression weights and inverse probability weights (IPW).
These findings suggest that applied researchers should interpret treatment effect estimates among under-represented subgroups with caution, given their sensitivity to modeling assumptions. As Parikh et al. (2025) emphasized, collecting more representative trial data is critical to credibly estimate treatment effects for this underrepresented subgroup in future medication for opioid use disorder studies.
Figure 5: Negative influence, defined as the contribution of negative weights in estimation, for different values of $\gamma$ and $\lambda$ along with implied linear regression weights and inverse probability weights (IPW).

### v2

@@ 3.1 Characterizing asymmetry-induced bias
Thus far we have presented a conservative nonparametric bound. We now provide a slightly refined characterization by noting that the extent of the bias induced by negative weights is driven by the asymmetry in $\mu$. We do so by considering the decomposition of the $\mu$ into its even and odd components, i.e., $\mu(x)=\mu_{e}(x)+\mu_{o}(x)$. By the definition of odd functions, we have $\mu_{o}(x)=\mu_{o}(-x)$, so we can bound the risk by bounding the worst-case risk of $\hat{\boldsymbol{w}}$ using the assumed Hölder constants $a$ and $\alpha$ and isolating the effect of the even component. The formal statement is given below in Proposition 3.1 the proof is given in Appendix C.
@@
Let $\hat{\mu}(x^{*})=\sum_{i=1}^{n}\hat{w}_{i}Y_{i}$ be the estimate of $\mu(x^{*})$ with weights estimated via Equation 3.2. Given $Y_{i}=\mu(X_{i})+\epsilon_{i}$ where $\epsilon_{i}$ are independent random variables with $\mathbb{E}[\epsilon_{i}]=0$ and finite second moment $\sigma^{2}=\mathbb{E}[\epsilon_{i}^{2}]$, and $\mu$ is Hölder continuous with constants $a$ and $\alpha$. If $\epsilon_{i}$ are sub-Gaussian [^2] We assume mean zero sub-Gaussian noise, analogous results can be obtained with this assumption replaced by bounded noise. with parameter $\sigma$, then with probability at least $1-\delta$,
@@
$B_{\text{even}}(x^{*})$ $=\big|\sum_{i=1}^{n}\hat{w}_{i}[\mu_{e}(X_{i})-\mu_{e}(x^{*})]+2\sum_{i=1}^{n}|\hat{w}_{i}|\mathbf{1}(\hat{w}_{i}<0)a\|X_{i}-x^{*}\|^{\alpha}\big|$
@@
The worst-case bound provided earlier is recovered if $\mu$ is an even function, i.e., contains no odd component.
One of the fundamental limitations of Proposition 3.1 is that it requires access to the separate even and odd functions constituting $\mu$ which are latent in practice. In our proposed estimator and empirical application we address this by preferring the conservative form which assumes the worst case form. For completeness we provide the following proposition that provides an empirical analog of Proposition 3.1. Intuitively, when the observations of $X$ are symmetric, i.e., for every $X_{i}$ $X_{i}$ is also in the dataset, then we can recover both the even and odd functions. In practice, because this symmetry is unlikely to hold we approximate it with a one-nearest neighbor and incorporate the induced uncertainty into the bound using the Hölder constants. The formal statement is below, the proof is deferred to Appendix C.
@@
Let $\hat{\mu}(x^{*})=\sum_{i=1}^{n}\hat{w}_{i}Y_{i}$ be the estimate of $\mu(x^{*})$ with weights estimated via Equation (5). Given $Y_{i}=\mu(X_{i})+\epsilon_{i}$ where $\epsilon_{i}$ are independent sub-Gaussian random variables with parameter $\sigma$, and $\mu$ is Hölder continuous with constants $a$ and $\alpha$. Let $I_{\text{paired}}=\{i:-X_{i}\in\{X_{1},\ldots,X_{n}\}\}$, $I_{\text{nn}}=\{1,\ldots,n\}\setminus I_{\text{paired}}$, and for each $i\in I_{\text{nn}}$, define $j^{*}(i)=\arg\min_{j\neq i}\|X_{j}-(-X_{i})\|$. Then with probability at least $1-\delta$, $|\mu(x^{*})-\hat{\mu}(x^{*})|$ is less than or equal to
$\left|\sum_{i\in I_{\text{paired}}}\hat{w}_{i}\frac{Y_{i}+Y_{-i}}{2}+\sum_{i\in I_{\text{nn}}}\hat{w}_{i}\frac{Y_{i}+Y_{j^{*}(i)}}{2}-\mu_{e}(x^{*})\sum_{i=1}^{n}\hat{w}_{i}\right|$
$+\sum_{i\in I_{\text{nn}}}|\hat{w}_{i}|a\|X_{j^{*}(i)}-(-X_{i})\|^{\alpha}$
$+2\sum_{i=1}^{n}|\hat{w}_{i}|\mathbf{1}(\hat{w}_{i}<0)a\|X_{i}-x^{*}\|^{\alpha}+\sigma\|\hat{w}\|_{2}\sqrt{2\log(6/\delta)}$
$+\frac{\sigma}{\sqrt{2}}(\left\|\hat{w}_{I_{\text{paired}}}\right\|_{2}+\left\|\hat{w}_{I_{\text{nn}}}\right\|_{2})\sqrt{2\log(6/\delta)}$
where $\mu_{e}(x^{*})$ is bounded using the closest observation $j^{*}=\arg\min_{j=1,\ldots,n}\|X_{j}-x^{*}\|$, $|\mu_{e}(x^{*})|$ less than or equal to
$\left|\frac{Y_{j^{*}}+Y_{j^{*}(j^{*})}}{2}\right|+\sigma\sqrt{2\log(6/\delta)}$
$+a\|X_{j^{*}(j^{*})}-(-X_{j^{*}})\|^{\alpha}+a\|X_{j^{*}}-x^{*}\|^{\alpha}$
Proposition 3.2 can be used as a companion to Chattopadhyay and Zubizarreta (2023), who propose computing the effective sample size of units with negative weight, $\frac{\sum_{i}\mathbf{1}[w_{i}<0]|w_{i}|}{\sum_{i}|w_{i}|}$, to indicate the extent to which the estimate relies on extrapolation and parametric assumptions – we refer to this as negative influence.
@@ 3.2 Learning Estimator
We now propose an estimator to learn weights $\boldsymbol{w}$} that directly control the error bound in Equation (2). To do so, we modify the standard balancing weights optimization problem in Equation (1) by using the Lagrangian form of the non-negative restriction, rather than the hard constraint. Thus, the combined estimator minimizes the error bound by controlling three terms: covariate imbalance, dispersion of the weights, and level of extrapolation:
$\hat{\mathbf{w}}\in\arg\min_{\mathbf{w}}\begin{pmatrix}\underbrace{\|\sum_{i}^{n}w_{i}{\boldsymbol{X}}_{i}-{\boldsymbol{x}}^{\star}\|^{2}_{p}}_{(a)}+\lambda\underbrace{\|\mathbf{w}\|^{2}_{2}}_{(b)}+\gamma\underbrace{\left\|\mathbbm{1}(w_{i}<0)|w_{i}|\left(\|{\boldsymbol{X}}_{i}\|^{\alpha}_{2}\right)\right\|_{p}}_{(c)}\end{pmatrix}$,
where
@@
The standard balancing weights problem in Equation (1) only focuses on a single bias-variance tradeoff: trading off covariate imbalance in the first term — which directly introduces bias — and the norm of the weights in the second term — which directly controls the estimator variance. By contrast, the new optimization problem (3.2) has a more elaborate bias-bias-variance trade-off. Allowing for negative weights introduces an additional trade-off between the first two terms and term (c): when the target unit lies outside the convex hull of the training points, controlling imbalance often requires some $w_{i}$ values to be negative, which also increases the norm $\|{\boldsymbol{w}}\|_{2}$. Allowing negative weights also necessitates reliance on parametric assumptions for extrapolation.
@@
Term (c) mitigates the risk of biased estimation by regulating the contribution of negative weights. For $\gamma=0$, Equation (3.2) recovers a standard, unconstrained balancing weights problem as in Equation (1). At the other extreme $\gamma\to\infty$ is equivalent to a hard non-negativity constraint. Increasing $\gamma$ constrains extrapolation, reducing bias due to possible violations of parametric assumptions and limiting $\|{\boldsymbol{w}}\|_{2}$, but worsening bias due to insufficient balance in term (a).
Hyperparameter choice. We argue that our proposed framework is best viewed as a form of sensitivity analysis, and encourage researchers to examine the full set of estimates spanned by the extrapolation hyperparameter $\gamma$. To supplement this, we propose a heuristic for choosing a default $\gamma^{*}$, starting with $\gamma=0$ (i.e., no penalty) and then gradually increasing $\gamma$ until all weights are non-negative i.e., $w\geq 0$. Alternatively, one can choose hyperparameter, in the spirit of Lepski’s method, by looking at when the change in the point estimate is sufficiently small, for a researcher-defined cutoff.
Finally, we provide the following proposition which specializes Proposition 3.2 for our proposed regularized estimator.
Proposition 3.3 (Regularized Bound).
Consider the regularized estimator,
$\hat{w}\in\arg\min_{w}$ $\left\|\sum_{i=1}^{n}w_{i}X_{i}-x^{*}\right\|_{2}^{2}+\lambda\|w\|_{2}^{2}+\gamma\sum_{i=1}^{n}\mathbf{1}(w_{i}<0)|w_{i}|\|X_{i}\|^{\alpha}$
and $\hat{\mu}(x^{*})=\sum_{i=1}^{n}\hat{w}_{i}Y_{i}$ where $Y_{i}=\mu(X_{i})+\epsilon_{i}$. Under the assumptions of Proposition 3.2 and $\|X_{i}\|\leq R$, with probability at least $1-\delta$:
$|\mu(x^{*})-\hat{\mu}(x^{*})|\leq$ $B_{\text{even}}(x^{*})+\sigma\left(\frac{\|x^{*}\|_{2}}{\sqrt{\lambda}}+\frac{\gamma R^{\alpha}\sqrt{n}}{\lambda}\right)\sqrt{2\log(2/\delta)}$
where $B_{\text{even}}(x^{*})$ is defined as in Proposition 3.2.
The proof is provided in the appendix, proceeding by providing a more explicit form for the weight norm in proposition 3.2. This formulation explicitly shows the trade-offs introduced by the nonnegative regularizer.
4 Generalizing Medication for Opioid Use Disorder Trial Evidence
@@
The Starting Treatment With Agonist Replacement Therapies (START) trial, initiated in 2006, was a multi-center study comparing buprenorphine versus methadone in treating opioid use disorder (Saxon et al., 2013; Hser et al., 2014). The trial enrolled 1,271 participants, who were randomized in a 2:1 ratio to receive either buprenorphine or methadone. Methadone was found to have higher rates of patient retention in treatment compared to buprenorphine (Hser et al., 2014). Our analysis focuses on the outcome of relapse to regular opioid use within 24 weeks of medication assignment, defined as non-study opioid use for four consecutive weeks or daily use for seven consecutive days.
@@
Parikh et al. (2025) identified that Hispanic women with a pre-treatment history of amphetamine and benzodiazepine use were underrepresented in the START trial relative to the target population, highlighting a practical violation of the positivity assumption. In this study, we estimate the target average treatment effects (TATE) for this underrepresented subgroup using our proposed framework alongside standard linear regression, gradient boosting regression (GBR), inverse probability weighting (IPW) and double machine learning estimators.
@@
The target sample is drawn from the 2015–2017 Treatment Episode Dataset - Admissions (TEDS-A), which includes data on individuals entering publicly funded substance use treatment programs across 48 states (excluding Oregon and Georgia) and the District of Columbia. Our analysis focuses on Hispanic women with a pre-treatment history of amphetamine and benzodiazepine use.
@@
Figure 1: Target Average Treatment Effects for the Target Sample for Hispanic Females who have a history of Amphetamize and Benzodiazepine use in TEDS-A population. Each hue corresponds to a value of $\lambda$ and the x-axis corresponds to different values of $\gamma$ (on log scale).
Figure 2: Balance between the trial and the target samples measured as the root mean squared error (RMSE) for different values of $\gamma$ and $\lambda$.
Figure 3: Negative influence, defined as the contribution of negative weights in estimation, for different values of $\gamma$ and $\lambda$.
We apply our proposed framework to address these issues, which regularizes extrapolation to mitigate reliance on extreme weights. By varying $\gamma$ from 0.01 to 10, we examine how treatment effect estimates shift with increasing regularization of negative weights. Without regularization, the point estimates converge to those from linear regression. However, as regularization intensifies, the estimates smoothly shift towards zero and occasionally change the sign from negative to positive for smaller values of $\lambda$. This sensitivity underscores the influence of assumptions on the point estimates. While increasing $\gamma$ reduces negative influence (Figure 3), it worsens covariate balance, as reflected in higher RMSE values (Figure 2). Thus, our framework highlights a trade-off between minimizing reliance on parametric assumptions and achieving optimal covariate balance. Our analysis highlights that these estimates are sensitivity to modeling assumptions. Thus, applied researchers should interpret treatment effect estimates among this under-represented subgroups with caution. As Parikh et al. (2025) emphasized, collecting more representative trial data is critical to credibly estimate treatment effects for this underrepresented subgroup.